# PlasmoData.jl – A Julia Framework for Modeling and Analyzing Complex Data as Graphs

### v1

@@ 2.3.3 Symmetric Matrices as Edge-Weighted Graphs
where $A$ is a symmetric matrix ($\mathcal{DG}$ is undirected) matching the connectivity defined by $\mathcal{E}$. The implementation for this representation is shown in Code Snippet 4. The symmetric matrix is passed to the function symmetric_matrix_to_graph (Line 9) which forms the edge weighted PlasmoData.DataGraph. The visualization of this is shown in Figure 6. The PlasmoData.DataGraph can also be visualized using the DataGraphPlots.jl, where we first set the node positions using set_circle_node_positions!, which provides the circular structure shown in Figure 6. The graph is visualized with the plot_graph function (Lines 15 - 22), where we color edges based on the edge weight through the line_z argument and the get_edge_data API from PlasmoData.jl (Line 21).
Code Snippet 4: Representing a symmetric matrix as a PlasmoData.DataGraph.
1using PlasmoData, Random, DataGraphPlots, LinearAlgebra
@@
4random_matrix = rand(6, 6)
5symmetric_matrix = (random_matrix .+ random_matrix) / 2
6symmetric_matrix[diagind(symmetric_matrix)] .= 1
7
8# Convert symmetric matrix to edge weighted graph
9symmetric_matrix_graph = symmetric_matrix_to_graph(symmetric_matrix)
10
11# Set node positions
12set_circle_node_positions!(symmetric_matrix_graph)
13
14# Plot the graph
15plot_graph(
@@
22)
Figure 6: Visualization of Code Snippet 4 where a PlasmoData.DataGraph is formed from a symmetric matrix.
@@ 3.1.1 Node and Edge Removal
The node- and edge-removal functions are simple operations but form the basis of more complex operations that manipulate the graph structure. The node- and edge-removal functions are implemented in PlasmoData as functions remove_node! and remove_edge!.
@@ 3.1.2 Graph Filtration
The above node- and edge-filtrations are implemented in PlasmoData.jl as shown in Code Snippet 5 and visualized in Figure 7. The graph defined on Line 4 can be filtered by calling the function filter_nodes (the software implementation of $f_{\mathcal{N}}$). This function takes arguments of the PlasmoData.DataGraph, a threshold value (often used in the filter function), the attribute name, and an optional keyword argument fn, which is the filter function used for the filtration. On Line 7, we filter out all nodes whose “weight” attribute is less than $0.7$ through the Base.isless function. On Lines 9 - 11, we define our own function for filtration called extreme_vals, which we pass to filter_nodes on Line 14 and which filters out all values that are not less than $0.2$ or greater than $0.8$. This highlights how user-defined functions can be used to filter out a graph. On Line 17, we add random weights to all the edges in the graph (note that add_edge_dataset! has a similar function as add_edge_data, but instead adds data of a single attribute to multiple edges, whereas the latter function only adds data to a single edge). Once data is defined on the edges, we can likewise perform filtrations based on the edge data, as shown on lines 20 and 23 where we now call filter_edges (the software implementation of $f_{\mathcal{E}}$). This function behaves similarly to filter_nodes, but now operates on the edge data rather than the node data. All arguments are the same for these functions, except that filter_edges requires an attribute defined on the edges rather than on the nodes. While these functions are shown for PlasmoData.DataGraphs, they also apply for PlasmoData.DataDiGraphs.
Code Snippet 5: Filtering a PlasmoData.DataGraph by node or edge data.
@@
2random_matrix = rand(12, 12)
3
4matrix_graph = matrix_to_graph(random_matrix; diagonal = true)
@@
6# Filter out nodes with weight =< 0.7
7filter_nodes_graph1 = filter_nodes(matrix_graph, 0.7, "weight", fn = Base.isless)
8
9function extreme_vals(a, b)
@@
11end
12
13# Filter out nodes with weight >= 0.2 and <= 0.8
14filter_nodes_graph2 = filter_nodes(matrix_graph, nothing, "weight", fn = extreme_vals)
15
16n_edges = length(matrix_graph.edges)
17add_edge_dataset!(matrix_graph, rand(n_edges), "weight")
18
19# Filter out edges with weight <= 0.8
20filter_edges_graph1 = filter_edges(matrix_graph, 0.8, "weight", fn = Base.isgreater)
21
22# Filter out edges with weight >= 0.5
23filter_edges_graph2 = filter_edges(matrix_graph, 0.5, "weight", fn = Base.isless)
Figure 7: Visualization of Code Snippet 5 where a PlasmoData.DataGraph is filtered based on node or edge data.
@@ 3.1.3 Graph Aggregation
Aggregation is a useful structural manipulation that can help reduce the dimensionality of the graph and enable scalable analysis. Graph aggregation is also often referred to as coarsening. For a DataGraph $\mathcal{DG}(\mathcal{N},\mathcal{E},A,\bm{d}^{\mathcal{N}},\bm{d}^{\mathcal{E}},\bm{d}^{\mathcal{G}},\mathcal{A}_{\mathcal{N}},\mathcal{A}_{\mathcal{E}},\mathcal{A}_{\mathcal{G}})$, aggregation combines a subset of nodes $\bar{\mathcal{N}}\subseteq\mathcal{N}$ into a new node, $n_{agg}$. This aggregation can also require aggregation of node and/or edge data.
Definition 3.5 (Aggregation Function $f_{agg}$).
@@
$\mathcal{DG}\left(\mathcal{N},\mathcal{E},A,\bm{d}^{\mathcal{N}},\bm{d}^{\mathcal{E}},\bm{d}^{\mathcal{G}},\mathcal{A}_{\mathcal{N}},\mathcal{A}_{\mathcal{E}},\mathcal{A}_{\mathcal{G}}\right)$ and a subset of nodes to be aggregated, $\bar{\mathcal{N}}\in\mathcal{N}$, the aggregation function collapses the node set into a single new node and aggregates node and edge data where applicable. The aggregation function $f_{agg}:(\mathbb{G},\mathbb{V},\mathbb{F},\mathbb{F})\rightarrow\mathbb{G}$ (where $\mathbb{V}$ is the set of all possible nodes and $\mathbb{F}$ is the set of all possible functions for aggregating sets of node or edge data) is defined as
@@
$\begin{split}f_{agg}(&\mathcal{DG},\mathcal{N}^{\prime},f^{\mathcal{N}}_{agg},f^{\mathcal{E}}_{agg})=\mathcal{DG}_{agg}\big{(}\mathcal{N}_{agg},\mathcal{E}_{agg},A_{agg}\bm{d}^{\mathcal{N}}_{agg},\bm{d}^{\mathcal{E}}_{agg},\bm{d}^{\mathcal{G}},\mathcal{A}_{\mathcal{N}},\mathcal{A}_{\mathcal{E}},\mathcal{A}_{\mathcal{G}}\big{)}\\ \textrm{where}&\;\mathcal{N}_{agg}=\{n_{agg}\}\cup\mathcal{N}\setminus\bar{\mathcal{N}}\\ &\;\bar{\mathcal{E}}_{agg}=\{(n_{i},n_{agg}):n_{i}\in\mathcal{N}\setminus\bar{\mathcal{N}},n_{j}\in\bar{\mathcal{N}},(n_{i},n_{j})\in\mathcal{E}\vee(n_{j},n_{i})\in\mathcal{E}\}\\ &\;\mathcal{E}_{agg}=\{(n_{i},n_{j}):n_{i}\not\in\bar{\mathcal{N}},n_{j}\not\in\bar{\mathcal{N}},(n_{i},n_{j})\in\mathcal{E}\}\cup\bar{\mathcal{E}}_{agg}\\ &\;\bm{d}^{\mathcal{N}}_{agg}(n)=\bm{d}^{\mathcal{N}}(n),\quad n\in\mathcal{N}\setminus\bar{\mathcal{N}}\\ &\;\bm{d}^{\mathcal{N}}_{agg}(n_{agg})=f^{\mathcal{N}}_{agg}(\bm{d}^{\mathcal{N}},\bar{\mathcal{N}})\\ &\;\bm{d}^{\mathcal{E}}_{agg}(e)=\bm{d}^{\mathcal{E}}(e),\quad e\in\{(n_{i},n_{j}):(n_{i},n_{j})\in\mathcal{E},n_{i}\not\in\bar{\mathcal{N}},n_{j}\not\in\bar{\mathcal{N}}\}\\ &\;\bm{d}^{\mathcal{E}}_{agg}((n_{i},n_{agg}))=f^{\mathcal{E}}_{agg}\big{(}\bm{d}^{\mathcal{E}},\{(n_{i},n_{j}):n_{j}\in\bar{\mathcal{N}},(n_{i},n_{j})\in\mathcal{E}\}\cup\\ &\;\qquad\qquad\qquad\qquad\qquad\qquad\qquad\{(n_{j},n_{i}):n_{j}\in\bar{\mathcal{N}},(n_{j},n_{i})\in\mathcal{E}\}\big{)}\quad(n_{i},n_{agg}),\in\bar{\mathcal{E}}_{agg}\end{split}$ (13)
@@
where $f^{\mathcal{N}}_{agg}$ is a function that aggregates node data in $\bm{d}^{\mathcal{N}}$ for nodes $\bar{\mathcal{N}}$ into node data for a single node ($n_{agg}$) and $f^{\mathcal{E}}_{agg}$ is a function that aggregates edge data in $\bm{d}^{\mathcal{E}}$ for the set of edges being passed to the function, and $A_{agg}$ is the adjacency matrix for edge set $\mathcal{E}_{agg}$.
@@
Note that the functions $f^{\mathcal{N}}_{agg}$ and $f^{\mathcal{E}}_{agg}$ can take many forms because of the variety of data that can be stored on the nodes and edges. In addition, the function for aggregating edge data, $f^{\mathcal{E}}_{agg}$, is only needed when there are multiple nodes in $\bar{\mathcal{N}}$ connected to some node in $\mathcal{N}\setminus\bar{\mathcal{N}}$; in this case, there are multiple edges that will be replaced by a single edge for undirected graphs. We also note that the aggregation function for directed graphs is different because the order of the edges will matter; consequently, rather than the function $f^{\mathcal{E}}_{agg}$ acting on the combined set of edges $\{(n_{i},n_{j}):n_{j}\in\bar{\mathcal{N}},(n_{i},n_{j})\in\mathcal{E}\}$ and $\{(n_{j},n_{i}):n_{j}\in\bar{\mathcal{N}},(n_{j},n_{i})\in\mathcal{E}\}$, the function could need to be applied to multiple, independent sets. In addition, the set $\bar{\mathcal{E}}$ would need to be expanded to consider the edges where $n_{agg}$ could be the source node (not just the destination) for an aggregated edge.
@@
PlasmoData.jl implements aggregation for graph objects, as shown in Code Snippet 6 and Figure 8. Here, a node set is defined in Line 7, and the node set is aggregated in line 10 using the function aggregate. The default for aggregating node data or edge data is to average the values of the data for each attribute, but one can also use, for instance, a $\max$ operator (as done in graph neural networks). The user can use these different functions by passing the function to the key word arguments node_fn and edge_fn.
Code Snippet 6: Aggregating a set of nodes in a PlasmoData.DataGraph
@@
2random_matrix = rand(12, 12)
3
4matrix_graph = matrix_to_graph(random_matrix; diagonal = true)
@@
6# Define nodes to be aggregated
7nodes_for_aggregation = [(3, 7), (3, 8), (3, 9), (4, 7), (4, 8)]
8
9# Aggregate nodes in the graph
10aggregated_graph = aggregate(matrix_graph, nodes_for_aggregation, "agg_node")
Figure 8: Visualization of Code Snippet 6 where a subset of nodes is aggregated into a single new node with the node weights on the aggregated node averaged to get a new node weight for the new node.

### v2

@@ 2.3.3 Symmetric Matrices as Edge-Weighted Graphs
PlasmoData.jl also facilitates this symmetric matrix graph representation. The implementation for this representation is shown in Code Snippet 4. The symmetric matrix is passed to the function symmetric_matrix_to_graph (Line 10) which forms the edge weighted PlasmoData.DataGraph. The visualization of this is shown in Figure 6. The PlasmoData.DataGraph can also be visualized using PlasmoDataPlots.jl, where we first set the node positions using set_circle_node_positions!, which provides the circular structure shown in Figure 6. The graph is visualized with the plot_graph function (Lines 16 - 23), where we color edges based on the edge weight through the line_z argument and the get_edge_data API from PlasmoData.jl (Line 22).
Code Snippet 4: Example of representing a symmetric matrix as a PlasmoData.DataGraph
1using PlasmoData, Random, PlasmoDataPlots, LinearAlgebra
@@
4Random.seed!(5)
5random_matrix = rand(6, 6)
6symmetric_matrix = (random_matrix .+ random_matrix) / 2
7symmetric_matrix[diagind(symmetric_matrix)] .= 1
8
9# Convert symmetric matrix to edge weighted graph
10symmetric_matrix_graph = symmetric_matrix_to_graph(symmetric_matrix)
11
12# Set node positions
13set_circle_node_positions!(symmetric_matrix_graph)
14
15# Plot the graph
16plot_graph(
@@
23)
Figure 6: Visualization of a graph representation of a symmetric matrix where the graph corresponds to the symmetric_matrix_graph PlasmoData.DataGraph object from Code Snippet 4. Values of matrix entries are stored as data on the edges of the graph
@@ 3.1.1 Node and Edge Removal
The node- and edge-removal functions are simple operations but form the basis of more complex operations that manipulate the graph structure. They also provide a basis for different analysis, such as analyzing the topology of a graph after the removal of a node or edge. The node- and edge-removal functions are implemented in PlasmoData as functions remove_node! and remove_edge!.
@@ 3.1.2 Graph Filtration
The above node- and edge-filtrations are implemented in PlasmoData.jl as shown in Code Snippet 5 and visualized in Figure 7. The graph defined on Line 6 can be filtered by calling the function filter_nodes (the software implementation of $f_{\mathcal{N}}$). This function takes arguments of the PlasmoData.DataGraph, a threshold value (often used in the filter function), the attribute name, and an optional keyword argument fn, which is the filter function used for the filtration. Note that all nodes where the filter function, fn, does not return true are filtered out. On Line 9, we filter out all nodes whose “weight” attribute is greater than $0.7$ through the Base.isless function. On Lines 11 - 13, we define our own function for filtration called extreme_vals, which we pass to filter_nodes on Line 16 and which filters out all values that are not less than $0.2$ or greater than $0.8$. This highlights how user-defined functions can be used to filter out a graph. On Line 19, we add random weights to all the edges in the graph (note that add_edge_dataset! has a similar function as add_edge_data, but instead adds data of a single attribute to multiple edges, whereas the latter function only adds data to a single edge). Once data is defined on the edges, we can likewise perform filtrations based on the edge data, as shown on lines 22 and 25 where we now call filter_edges (the software implementation of $f_{\mathcal{E}}$). This function behaves similarly to filter_nodes, but now operates on the edge data rather than the node data. All arguments are the same for these functions, except that filter_edges requires an attribute defined on the edges rather than on the nodes. While these functions are shown for PlasmoData.DataGraphs, they also apply for PlasmoData.DataDiGraphs.
Code Snippet 5: Example of filtering a PlasmoData.DataGraph by node or edge data
@@
2
3Random.seed!(15)
4random_matrix = rand(12, 12)
@@
6matrix_graph = matrix_to_graph(random_matrix; diagonal = true)
7
8# Filter out/remove nodes with weight >= 0.7
9filter_nodes_graph1 = filter_nodes(matrix_graph, 0.7, "weight", fn = Base.isless)
10
11function extreme_vals(a, b)
@@
13end
14
15# Filter out/remove nodes with weight >= 0.2 and <= 0.8
16filter_nodes_graph2 = filter_nodes(matrix_graph, nothing, "weight", fn = extreme_vals)
17
18n_edges = length(matrix_graph.edges)
19add_edge_dataset!(matrix_graph, rand(n_edges), "weight")
20
21# Filter out/remove edges with weight <= 0.8
22filter_edges_graph1 = filter_edges(matrix_graph, 0.8, "weight", fn = Base.isgreater)
23
24# Filter out/remove edges with weight >= 0.5
25filter_edges_graph2 = filter_edges(matrix_graph, 0.5, "weight", fn = Base.isless)
Figure 7: Visualization of different examples of node or edge filtration associated with Code Snippet 5, where different logic sets are used for removing nodes or edges of the graph
@@ 3.1.3 Graph Aggregation
Aggregation is a useful structural manipulation that can help reduce the dimensionality of the graph and enable scalable analysis. Graph aggregation is also often referred to as coarsening. For a DataGraph $\mathcal{DG}(\mathcal{N},\mathcal{E},A,\bm{d}^{\mathcal{N}},\bm{d}^{\mathcal{E}},\bm{d}^{\mathcal{G}},\mathcal{A}_{\mathcal{N}},\mathcal{A}_{\mathcal{E}},\mathcal{A}_{\mathcal{G}})$, aggregation combines a subset of nodes $\bar{\mathcal{N}}\subseteq\mathcal{N}$ into a new node, $n_{agg}$. This aggregation can also require aggregation of node and/or edge data. We formalize the aggregation process for DataGraphs by defining the following:
Definition 3.5 (Aggregation Function $f_{+}$).
@@
$\mathcal{DG}\left(\mathcal{N},\mathcal{E},A,\bm{d}^{\mathcal{N}},\bm{d}^{\mathcal{E}},\bm{d}^{\mathcal{G}},\mathcal{A}_{\mathcal{N}},\mathcal{A}_{\mathcal{E}},\mathcal{A}_{\mathcal{G}}\right)$ and a subset of nodes to be aggregated, $\bar{\mathcal{N}}\in\mathcal{N}$, the aggregation function collapses the node set into a single new node and aggregates node and edge data where applicable. The aggregation function $f_{+}:(\mathbb{G},\mathbb{V},\mathbb{F},\mathbb{F})\rightarrow\mathbb{G}$ (where $\mathbb{V}$ is the set of all possible nodes and $\mathbb{F}$ is the set of all possible functions for aggregating sets of node or edge data) is defined as
@@
$\begin{split}f_{+}(&\mathcal{DG},\mathcal{N}^{\prime},f^{\mathcal{N}}_{+},f^{\mathcal{E}}_{+})=\mathcal{DG}_{agg}\big{(}\mathcal{N}_{agg},\mathcal{E}_{agg},A_{agg},\bm{d}^{\mathcal{N}}_{agg},\bm{d}^{\mathcal{E}}_{agg},\bm{d}^{\mathcal{G}},\mathcal{A}_{\mathcal{N}},\mathcal{A}_{\mathcal{E}},\mathcal{A}_{\mathcal{G}}\big{)}\\ \textrm{where}&\;\mathcal{N}_{agg}=\{n_{agg}\}\cup\mathcal{N}\setminus\bar{\mathcal{N}}\\ &\;\bar{\mathcal{E}}_{agg}=\{(n_{i},n_{agg}):n_{i}\in\mathcal{N}\setminus\bar{\mathcal{N}},n_{j}\in\bar{\mathcal{N}},(n_{i},n_{j})\in\mathcal{E}\vee(n_{j},n_{i})\in\mathcal{E}\}\\ &\;\mathcal{E}_{agg}=\{(n_{i},n_{j}):n_{i}\not\in\bar{\mathcal{N}},n_{j}\not\in\bar{\mathcal{N}},(n_{i},n_{j})\in\mathcal{E}\}\cup\bar{\mathcal{E}}_{agg}\\ &\;\bm{d}^{\mathcal{N}}_{agg}(n)=\bm{d}^{\mathcal{N}}(n),\quad n\in\mathcal{N}\setminus\bar{\mathcal{N}}\\ &\;\bm{d}^{\mathcal{N}}_{agg}(n_{agg})=f^{\mathcal{N}}_{+}(\bm{d}^{\mathcal{N}},\bar{\mathcal{N}})\\ &\;\bm{d}^{\mathcal{E}}_{agg}(e)=\bm{d}^{\mathcal{E}}(e),\quad e\in\{(n_{i},n_{j}):(n_{i},n_{j})\in\mathcal{E},n_{i}\not\in\bar{\mathcal{N}},n_{j}\not\in\bar{\mathcal{N}}\}\\ &\;\bm{d}^{\mathcal{E}}_{agg}((n_{i},n_{agg}))=f^{\mathcal{E}}_{+}\big{(}\bm{d}^{\mathcal{E}},\{(n_{i},n_{j}):n_{j}\in\bar{\mathcal{N}},(n_{i},n_{j})\in\mathcal{E}\}\cup\\ &\;\qquad\qquad\qquad\qquad\qquad\qquad\qquad\{(n_{j},n_{i}):n_{j}\in\bar{\mathcal{N}},(n_{j},n_{i})\in\mathcal{E}\}\big{)}\quad(n_{i},n_{agg}),\in\bar{\mathcal{E}}_{agg}\end{split}$ (13)
@@
where $f^{\mathcal{N}}_{+}$ is a function that aggregates node data in $\bm{d}^{\mathcal{N}}$ for nodes $\bar{\mathcal{N}}$ into node data for a single node ($n_{agg}$) and $f^{\mathcal{E}}_{+}$ is a function that aggregates edge data in $\bm{d}^{\mathcal{E}}$ for the set of edges being passed to the function, and $A_{agg}$ is the adjacency matrix for edge set $\mathcal{E}_{agg}$.
@@
In the above definition, we introduce two new functions, $f^{\mathcal{N}}_{+}$ and $f^{\mathcal{E}}_{+}$. These functions can take many forms because of the variety of data that can be stored on the nodes and edges. In addition, the function for aggregating edge data, $f^{\mathcal{E}}_{+}$, is only needed when there are multiple nodes in $\bar{\mathcal{N}}$ connected to some node in $\mathcal{N}\setminus\bar{\mathcal{N}}$; in this case, there are multiple edges that will be replaced by a single edge for undirected graphs. We also note that the aggregation function for directed graphs is different because the order of the edges will matter; consequently, rather than the function $f^{\mathcal{E}}_{+}$ acting on the combined set of edges $\{(n_{i},n_{j}):n_{j}\in\bar{\mathcal{N}},(n_{i},n_{j})\in\mathcal{E}\}$ and $\{(n_{j},n_{i}):n_{j}\in\bar{\mathcal{N}},(n_{j},n_{i})\in\mathcal{E}\}$, the function could need to be applied to multiple, independent sets. In addition, the set $\bar{\mathcal{E}}$ would need to be expanded to consider the edges where $n_{agg}$ could be the source node (not just the destination) for an aggregated edge. For both directed and undirected graphs, $f^{\mathcal{N}}_{+}$ and $f^{\mathcal{E}}_{+}$ must be permutation invariant.
@@
PlasmoData.jl implements aggregation for graph objects, as shown in Code Snippet 6 and Figure 8. Here, a node set is defined in Line 9, and the node set is aggregated in line 12 using the function aggregate. The default for aggregating node data or edge data is to average the values of the data for each attribute, but one can also use, for instance, a $\max$ operator (as done in graph neural networks). The user can use these different functions by passing the function to the key word arguments node_fn and edge_fn.
Code Snippet 6: Example of aggregating a set of nodes in a PlasmoData.DataGraph
@@
2
3Random.seed!(15)
4random_matrix = rand(12, 12)
@@
6matrix_graph = matrix_to_graph(random_matrix; diagonal = true)
7
8# Define nodes to be aggregated
9nodes_for_aggregation = [(3, 7), (3, 8), (3, 9), (4, 7), (4, 8)]
10
11# Aggregate nodes in the graph
12aggregated_graph = aggregate(matrix_graph, nodes_for_aggregation, "agg_node")
Figure 8: Visualization associated with Code Snippet 6 where a subset of nodes is aggregated into a single new node with the node weights on the aggregated node averaged to get a new node weight for the new node
There is an extension of (13) that could be considered. Under (13), any data on the edges between nodes in $\bar{\mathcal{N}}$ is lost, but such data could be important in different applications. We could extend (13) by defining a function $f_{+}^{\mathcal{E}\rightarrow\mathcal{N}}$ that reduces $\bm{d}^{\mathcal{E}}(e),e\in\{(n_{i},n_{j}):(n_{i},n_{j})\in\mathcal{E},n_{i}\in\bar{\mathcal{N}},n_{j}\in\bar{\mathcal{N}}\}$ into new data on $n_{agg}$ for an additional set of attributes, $\mathcal{A}_{\mathcal{E}\rightarrow\mathcal{N}}$. Such attributes would not be restricted to be in either $\mathcal{A}_{\mathcal{N}}$ or $\mathcal{A}_{\mathcal{E}}$. This would result in a slightly different DataGraph such that $\mathcal{DG}_{agg}\big{(}\mathcal{N}_{agg},\mathcal{E}_{agg},A_{agg},\bm{d}^{\mathcal{N}}_{agg},\bm{d}^{\mathcal{E}}_{agg},\bm{d}^{\mathcal{G}},\mathcal{A}_{\mathcal{N}}\cup\mathcal{A}_{\mathcal{E}\rightarrow\mathcal{N}},\mathcal{A}_{\mathcal{E}},\mathcal{A}_{\mathcal{G}}\big{)}$, where $\bm{d}^{\mathcal{N}}_{agg}$ now contains additional data for $\mathcal{A}_{\mathcal{E}\rightarrow\mathcal{N}}$. PlasmoData.jl does not (yet) support this extension, so we have restricted (13) to follow the implementation of PlasmoData.jl, but this could be an area of future work.
note that aggregation here refers to reducing the number of nodes in the graph and requires a defined set of nodes to be aggregated. The definition for fN+ and fE+ must be permutation invariant; in addition,